On February 1, 2021, Arrow Construction Company entered into a three-year construction contract to build a bridge for a price+ = % complete to date 2021 $ Percentages of completion Choose numerator Choose denominator Actual costs to date Estimated toReg 1 Req 2A Req 2B Req ЗА Req 3B Prepare journal entries for 2021 to record the transactions described (credit various accoReq 1 Reg 2A Req 2B Req ЗА Req 3B Prepare journal entries for 2022 to record the transactions described (credit various accoReg 1 Req 2A Req 2B Req ЗА Req 3B Prepare a partial balance sheet to show the presentation of the project as of December 31,Req 1 Req 2A Req 2B Req ЗА Reg 3B Prepare a partial balance sheet to show the presentation of the project as of December 31,
Answer #1

1. Computation of amount of revenue and gross profit using percentage of completion method.

Particulars 2021 2022 2023
Contract price 8,300,000 8,300,000 8,300,000
Cost incurred during the year 2,100,000 2,600,000 3,900,000
Estimated cost to complete (A) 4,100,000 3,750,000 -
Cost incurred up-to date (Sum of total costs incurred during the year) B 2,100,000 4,700,000 8,600,000
Estimated total cost (A+B) 6,200,000 8,450,000 8,600,000
% of completion (Cost incurred up-to date /total cost ) 33.87% 55.62% 100%
Revenue to be recognized

(33.87%*8,300,000)

=2,811,210

(55.62%*8,300,000-2,811,210)

=1,805,250

(100%*8,300,000-2,811,210-1,805,250)

=3,683,540

Gross profit /(loss) = Revenue to be recognized -cost incurred during the year 711,210 (794,750) (216,460)
Bills during the year 2,600,000 2,850,000
Collections during the year 2,350,000 2,575,000

2(a)&(b) 2021. 2022  

Journal entries Debit Credit Debit Credit
Costs incurred  

CIP

To Cash

2100,000 2100,000 2600,000 2,600,000
Billing

AR

To billing on CIP

2600,000 2600,000 2850,000 2,850,000
Receiving payment

cash

To AR

2,350,000 2,350,000 2,575,000 2,575,000
Period end adjustment for revenue (GP)

Construction expense

CIP(WIP)

To Construction Revenue

2100,000

711,210

2,811,210

Period end adjustment for revenue (GL)

construction expense

To CIP(WIP)

Construction revenue  

2,600,000

794,750

1,805,250

3.(a) &(b) partial balancesheet

Particulars 2021 2022
CA
CIP(WIP) 2100,000+711,210=2,811,210 2,811,210+2,600,000-794,750=4,616,460
Accounts receivable 2,600,000-2,350,000=250,000 2,850,000-2,575,000+250000=525,000
Total 3,061,210 5,141,460
CL
Billing on CIP 2,600,000 5,450,000(2600,000+2850,000)